Eight years ago, Hold was 7 times older than her son. Today, she is exactly 3 times as old as her son.
How old are both Mrs. Hold and her son today?

Answer by Cromlix(4381)

Hold was 7x older than her son.
H = 7S
Eight years later she is exactly
3 x as old as her son.
H + 8 = 3(S + 8)
Replace H with 7S from first equation
7S + 8 = 3(S + 8)
7S + 8 = 3S + 24
4S = 16
S = 4
Ages today.
Son = 12 years old
Mrs. Hold = 36 years old.
Hope this helps.
:-)